# Towards a Unified Framework for Evaluating Explanations

### v1

@@ 1 Introduction
The growing awareness in educational data mining (EDM) of a need for more explainable AI (XAI) has led to the increasing adoption of interpretability methods. Such methods are continually being developed and refined within research communities such as machine learning (ML) and human-computer interaction (HCI). However, the critical task of evaluating the efficacy of the explanations created has not been sufficiently explored. Tellingly, a systematic review of explainable student performance prediction models did not find a single study that evaluated the explanations it produced [1]. Furthermore, a standardized framework for conducting such evaluations is still lacking [15].
@@ 4 Evaluation criteria
Figure 1: Evaluation criteria framework. Edges depict the direction of dependence (A -> B = A depends on B).
@@ 6 Evaluation methods
  • Functionally-grounded evaluation, which does not involve humans but rather uses quantifiable properties of explanations as a proxy for interpretability.
An example of application grounded evaluation in education is the way learning dashboards are sometimes evaluated by how well they help instructors understand and provide help to students [21, 25], or work on open learner models (OLMs) that provide students with explanations of a model’s estimates of their understanding [7]. This evaluation method can be used to effectively measure explanation intelligibility (and, by extension, plausibility), but it does not directly tackle the question of explanation faithfulness.
Functionally-grounded evaluation, being the least direct category, makes it difficult to make any claims about either intelligibility or faithfulness. It allows for a proxy measurement of intelligibility by considering properties such as model sparsity or explanation simplicity [17], but it does not capture the specific needs of any end-user. While it can be helpful to consider potential target users while conducting this type of evaluation—ideally realistic stakeholders in education—the results are generally context-agnostic and therefore may lack real-world validity. Stability is perhaps the only criterion that can effectively be evaluated using functionally-grounded evaluation. This method may be most appropriate for preliminary studies in an area without much prior research.

### v2

@@ 1 Introduction
The growing awareness in educational data mining (EDM) of a need for more explainable AI (XAI) has led to the increasing discussion and adoption of interpretability methods [12]. Such methods are continually being developed and refined within research communities such as machine learning (ML) and human-computer interaction (HCI). However, the critical task of evaluating the efficacy of the explanations created has not been sufficiently explored. Tellingly, a systematic review of explainable student performance prediction models did not find a single study that evaluated the explanations it produced [1]. Furthermore, a standardized framework for conducting such evaluations is still lacking [16].
@@ 4 Evaluation criteria
Figure 1: Evaluation criteria framework. Edges depict the direction of dependence (A -> B = A is a prerequisite of B).
@@ 6 Evaluation methods
  • Functionally grounded evaluation, which does not involve humans but rather uses quantifiable properties of explanations as a proxy for interpretability.
An example of application-grounded evaluation in education is the way learning dashboards are sometimes evaluated by how well they help instructors understand and provide help to students [22, 26], or work on open learner models (OLMs) that provide students with explanations of a model’s estimates of their understanding [7]. This evaluation method can be used to effectively measure explanation intelligibility (and, by extension, plausibility), but it does not directly tackle the question of explanation faithfulness.
Functionally grounded evaluation, being the least direct category, makes it difficult to make any claims about either intelligibility or faithfulness. It allows for a proxy measurement of intelligibility by considering properties such as model sparsity or explanation simplicity [18], but it does not capture the specific needs of any end-user. While it can be helpful to consider potential target users while conducting this type of evaluation—ideally realistic stakeholders in education—the results are generally context-agnostic and therefore may lack real-world validity. Stability is perhaps the only criterion that can effectively be evaluated using functionally grounded evaluation. This method may be most appropriate for preliminary studies in an area without much prior research.